Two planets are separated in space by some distance d, each orbiting around their center of mass in the middle of them. They both have the same mass m = 4.81 × 10^20 kg and are rotating with a constant ω = 1.25 × 10^−10 rad/s. How far apart are they (d)?

The answer is 1.6 x 10^10 meters.
Answer #1

The two planets are orbiting around each other about their centre of mass. Let the distances between the planets from the centre of mass of their system be d1 and d2.

The distance between the two planets would now be equal to d = d1 + d2

Let the centre of mass be at x = 0

therefore the centre of mass equation will be:

\vec{X} = \frac{M_1d_1 + M_2(-d_2)}{M}

0 = \frac{M_1d_1 - M_2d_2}{M}

therefore, M1d1 = M2d2

and d = d1 + d2

and so d = d1 + (d1M1/M2)

writing it in terms of d2 we get,

d_2 = \frac{M_1}{M_1 +M_2}d.......................(1)

and we also know that for a system like this, the Gravitational Force provided by M1 gives the necessary centripetal force to M2 to remain in the orbit.

F_{grav} = GM_1M_2/d^2............................(2)

F_{centripetal} = M_2\omega^2 d_2..........................(3)

Notice the distance in the centripetal force is not d since M2 revolves around the centre of mass, not M1.

therefore,

so equating equating (1) and (2), we get:

GM_1M_2/d^2 = M_2\omega^2 d_2

or \frac{GM_1M_2}{d^2M_2d_2} = \omega^2

substitute the value of d2 from equation (1)

\frac{GM_1M_2}{d^2M_2 \frac{M_1d}{M_1 +M_2}} = \omega^2

\frac{G(M_1 +M_2)}{d^3} = \omega^2

or d^3 = \frac{G(M_1 +M_2)} {\omega^2}

but M1 = M2

therefore, d^3 = \frac{G2M} {\omega^2}

plug in the value of G = 6.67 x 10-11 m3kg-1s-2 , M = 4.81 x 1020kg and angular velocity = 1.25 x 10-10 rad/s to find the value of d.

d = 1.601 x 1010 m